# Deep Variational Network Toward Blind Image Restoration

### v5

@@ 2 Related Works
2.1 Model-based Methods
@@ 2.2 DL-based Methods
Inspired by the development of generative adversarial network (GAN) [47], some DL-based methods also followed the research line of noise modeling introduced in Sec. 2.1. Typically, Chen et al. [48] proposed a noise generator to simulate the real noise under the adversarial training mechanism, and Kim et al. [49] further introduced some camera settings (e.g., ISO level and shutter speed) into the generator as extra guidance. More recently, Yue et al. [24] proposed a dual adversarial loss to implement the noise removal and noise generation tasks into one unique Bayesian framework. Different from these GAN-based implicit noise modeling manners, this study adopts a more powerful and flexible non-i.i.d. Gaussian distribution to explicitly model the image noise, which avoids the instability in training GAN.
2.3 Some Relevant Explorations
@@
3 The Proposed Method
3.1 Basic Settings
@@
3.2 Model Formulation for Denoising
@@
where $\mathcal{N}(\cdot|\mu,\sigma^{2})$ denotes the Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$, $\bm{z}$ denotes the latent clean image, $d$ is the number of pixels in the noisy image. Notably, different from the commonly used i.i.d. Gaussian/Laplacian assumption, we model the image noise as pixel-wise non-i.i.d. Gaussian distribution in Eq. (2). Such an non-i.i.d. noise assumption largely increases the degrees of freedom of the noise distribution and is thus expected to better fit complicated real noise as illustrated in Sec. 5.3.1.
@@
Actually, $\xi_{i}$ in Eq. (5) provides an estimation for the variance $\sigma_{i}^{2}$. It is calculated using a Gaussian filter in the $p\times p$ window centered at the $i$-th pixel. The elaborate design on the shape and scale parameters (i.e., $\alpha_{0}-1$ and $\alpha_{0}\xi_{i}$) in Eq. (4) guarantees that the mode of this prior distribution is $\xi_{i}$ exactly. The hyper-parameter $\alpha_{0}$ controls the strength of this prior distribution, and it is set as $\nicefrac{{p^{2}}}{{2}}$ in our method following [67]. More explanations on this hyper-parameter can be found in Appendix .1.
@@
3.3 Extension to Blind Super-resolution
@@
4 Stochastic Variational Inference
@@
4.1 Form of Variational Posterior
@@
where $\beta_{i}(\bm{y};W_{S})$ is a mapping function parameterized as a DNN named as sigma network (SNet) with parameters $W_{S}$. It aims to predict the scale parameter of $q(\bm{\sigma}^{2}|\bm{y})$ directly from the corrupted image $\bm{y}$. As for the shape parameter of $q(\bm{\sigma}^{2}|\bm{y})$, we simply fix it as the same to the prior distribution, i.e., $\alpha_{0}-1$. Different from the strategy of setting them both as learnable parameters in our previous version [27], such a modification largely simplifies the evidence lower bound of Sec. 4.2 and also makes our algorithm more stable during training. Similarly, we formulate $q(\bm{\varLambda}|\bm{y})$ as:
@@
4.2 Evidence Lower Bound
@@ 4.2.1 Derivation of ELBO
4.2.2 Discussion
Let us delve further into the ELBO loss of Eq. (18). The first term of this formulation presents the log-likelihood of the observed low-resolution image, enforcing that the recovered high-resolution image can be mapped back to the low-resolution one via the estimated degradation model. The other three terms correspond to the constraints for the estimated high-resolution image $\bm{z}$, the noise variance $\bm{\sigma}^{2}$, and the kernel parameters $\bm{\varLambda}$, respectively, as shown in Eqs. (21)-(23). Particularly noteworthy is the constraint imposed on $\bm{z}$ in Eq. (21), which takes the form of an MSE endowed with a weight factor of $\frac{1}{2\varepsilon_{0}^{2}}$. In other words, our designed ELBO loss degenerates into the commonly used MSE by setting the hyper-parameter $\varepsilon_{0}$ to a small value close to zero. A more thorough experimental comparison with the MSE baseline can be found in Sec. 5.4.
@@ 4.2.2 Discussion
4.3 Network Structure and Learning
@@
It should be noted that this work does not aim to design more effective network architectures to surpass current SotA methods but to devise a probabilistic framework based on the deep variational inference for blind IR. Therefore, we simply employ the commonly used networks in low-level vision as our backbones. To better verify the generality of the proposed framework, we also test its effectiveness on other popular network architectures in Sec. 5.4.3.
5 Experimental Results
@@
5.1 Image Denoising Experiments
5.1.1 Synthetic Non-I.I.D. Gaussian Noise Removal
@@ 5.1 Image Denoising Experiments
5.1.2 Real-world Noise Removal
@@
5.2 Image Super-resolution Experiments
@@
5.2.1 Results on Synthetic Data
@@
Table V lists the comparison results of different methods under a scale factor of 4, and more results under scale factors of 2 and 3 are put into Appendix .3. From Table V, it can be seen that the proposed VIRNet achieves the best performance among blind methods in all cases. Especially, compared with the non-blind methods, VIRNet is still able to obtain slightly better or at least comparable results even though they make use of the groud truth information of blur kernel and noise level. This indicates the effectiveness of the proposed blind framework which is capable of handling the tasks of degradation estimation and image restoration simultaneously. Further, taking the model profiles into consideration, the superiority of VIRNet is more evident. Specifically, VIRNet has fewer number of parameters, fewer FLOPs, and faster speed than both the SotA blind method DASR [70] and the non-blind method USRNet [56].
@@
5.2.2 Results on Real Data
@@
5.2.3 Discussion on Blur Kernel
@@
5.3 Degradation Estimation Experiments
@@ 5.3.1 Noise Estimation
Non-I.I.D. Gaussian Noise. In Sec. 5.1.1, we adopt three groups of noise variance maps (see Fig. 3 (b1-d1)) to synthesize the testing data, to evaluate the performance of VIRNet under the non-i.i.d. Gaussian noise. Correspondingly, Fig. 3 (b2-d2) further displays the variance maps predicted by VIRNet for easy visualization. It can be seen that these predicted variance maps have very similar spatial variation with the ground truth ones, which are expected to facilitate the subsequent denoising task or other non-blind denoising methods. To justify this point, we also apply these predicted variance maps in FFDNet [20] to test its performance under the non-i.i.d. Gaussian noise, and the quantitative comparisons are listed in Table VIII. One can see that $\text{FFDNet}_{\text{VIR}}$ and $\text{FFDNet}_{\text{GT}}$ have very similar performance, and the performance difference between them is less than 0.04dB PSNR. This indicates that VIRNet is able to effectively handle such complicated noise distribution.
@@ 5.3.2 Kernel Estimation
5.4 Ablation Study
5.4.1 Posteriori Factorization
When designing the variational inference algorithm in Sec. 4, we factorize the variational distribution $q(\bm{z},\bm{\sigma}^{2},\bm{\varLambda}|\bm{y})$ into a conditional format of Eq. (12), which fundamentally induces the cascaded inference framework in Fig. 2. In fact, different factorized assumptions on $q(\bm{z},\bm{\sigma}^{2},\bm{\varLambda}|\bm{y})$ will lead to different designs on the inference framework. For example, the following unconditional factorization
@@ 5.4.1 Posteriori Factorization
5.4.2 Hyper-parameter analysis
@@
Hyper-parameter $\varepsilon_{0}^{2}$. As discussed in Sec. 4.2.2, Our ELBO loss degenerates to MSE when $\varepsilon_{0}$ is set as an extremely small value. We thus trained the RNet separately under MSE as a baseline for explicit comparison. The evaluative performance of VIRNet under different configurations of $\varepsilon_{0}$ is summarized in Table XII. This table unveils the following observations: 1) Too large setting on $\varepsilon_{0}^{2}$ (e.g., 1e-4) yields a relatively inferior performance of VIRNet, attributed to the attenuation of the prior constraint on $\bm{z}$; 2) our VIRNet performs well as long as $\varepsilon_{0}^{2}$ resides within the range of [1e-5, 1e-7], and deteriorates obviously when $\varepsilon_{0}$ is smaller beyond the threshold of 1e-7; 3) VIRNet surpasses the MSE baseline by 0.12dB in terms of PSNR, indicating the pivotal role of noise modeling in our proposed methodology. Thus, we suggest to set $\varepsilon_{0}^{2}$ as 1e-6 in this work.

### v6

@@ II Related Works
II-A Model-based Methods
@@ II-B DL-based Methods
Inspired by the development of generative adversarial network (GAN) [47], some DL-based methods also followed the research line of noise modeling introduced in Sec. II-A. Typically, Chen et al. [48] proposed a noise generator to simulate the real noise under the adversarial training mechanism, and Kim et al. [49] further introduced some camera settings (e.g., ISO level and shutter speed) into the generator as extra guidance. More recently, Yue et al. [24] proposed a dual adversarial loss to implement the noise removal and noise generation tasks into one unique Bayesian framework. Different from these GAN-based implicit noise modeling manners, this study adopts a more powerful and flexible non-i.i.d. Gaussian distribution to explicitly model the image noise, which avoids the instability in training GAN.
II-C Some Relevant Explorations
@@
III The Proposed Method
III-A Basic Settings
@@
III-B Model Formulation for Denoising
@@
where $\mathcal{N}(\cdot|\mu,\sigma^{2})$ denotes the Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$, $\bm{z}$ denotes the latent clean image, $d$ is the number of pixels in the noisy image. Notably, different from the commonly used i.i.d. Gaussian/Laplacian assumption, we model the image noise as pixel-wise non-i.i.d. Gaussian distribution in Eq. (2). Such an non-i.i.d. noise assumption largely increases the degrees of freedom of the noise distribution and is thus expected to better fit complicated real noise as illustrated in Sec. V-C1.
@@
Actually, $\xi_{i}$ in Eq. (5) provides an estimation for the variance $\sigma_{i}^{2}$. It is calculated using a Gaussian filter in the $p\times p$ window centered at the $i$-th pixel. The elaborate design on the shape and scale parameters (i.e., $\alpha_{0}-1$ and $\alpha_{0}\xi_{i}$) in Eq. (4) guarantees that the mode of this prior distribution is $\xi_{i}$ exactly. The hyper-parameter $\alpha_{0}$ controls the strength of this prior distribution, and it is set as $\nicefrac{{p^{2}}}{{2}}$ in our method following [67]. More explanations on this hyper-parameter can be found in Appendix -A.
@@
III-C Extension to Blind Super-resolution
@@
IV Stochastic Variational Inference
@@
IV-A Form of Variational Posterior
@@
where $\beta_{i}(\bm{y};W_{S})$ is a mapping function parameterized as a DNN named as sigma network (SNet) with parameters $W_{S}$. It aims to predict the scale parameter of $q(\bm{\sigma}^{2}|\bm{y})$ directly from the corrupted image $\bm{y}$. As for the shape parameter of $q(\bm{\sigma}^{2}|\bm{y})$, we simply fix it as the same to the prior distribution, i.e., $\alpha_{0}-1$. Different from the strategy of setting them both as learnable parameters in our previous version [27], such a modification largely simplifies the evidence lower bound of Sec. IV-B and also makes our algorithm more stable during training. Similarly, we formulate $q(\bm{\varLambda}|\bm{y})$ as:
@@
IV-B Evidence Lower Bound
@@ IV-B1 Derivation of ELBO
IV-B2 Discussion
Let us delve further into the ELBO loss of Eq. (18). The first term of this formulation presents the log-likelihood of the observed low-resolution image, enforcing that the recovered high-resolution image can be mapped back to the low-resolution one via the estimated degradation model. The other three terms correspond to the constraints for the estimated high-resolution image $\bm{z}$, the noise variance $\bm{\sigma}^{2}$, and the kernel parameters $\bm{\varLambda}$, respectively, as shown in Eqs. (21)-(23). Particularly noteworthy is the constraint imposed on $\bm{z}$ in Eq. (21), which takes the form of an MSE endowed with a weight factor of $\frac{1}{2\varepsilon_{0}^{2}}$. In other words, our designed ELBO loss degenerates into the commonly used MSE by setting the hyper-parameter $\varepsilon_{0}$ to a small value close to zero. A more thorough experimental comparison with the MSE baseline can be found in Sec. V-D.
@@ IV-B2 Discussion
IV-C Network Structure and Learning
@@
It should be noted that this work does not aim to design more effective network architectures to surpass current SotA methods but to devise a probabilistic framework based on the deep variational inference for blind IR. Therefore, we simply employ the commonly used networks in low-level vision as our backbones. To better verify the generality of the proposed framework, we also test its effectiveness on other popular network architectures in Sec. V-D3.
V Experimental Results
@@
V-A Image Denoising Experiments
V-A1 Synthetic Non-I.I.D. Gaussian Noise Removal
@@ V-A Image Denoising Experiments
V-A2 Real-world Noise Removal
@@
V-B Image Super-resolution Experiments
@@
V-B1 Results on Synthetic Data
@@
Table V lists the comparison results of different methods under a scale factor of 4, and more results under scale factors of 2 and 3 are put into Appendix -C. From Table V, it can be seen that the proposed VIRNet achieves the best performance among blind methods in all cases. Especially, compared with the non-blind methods, VIRNet is still able to obtain slightly better or at least comparable results even though they make use of the groud truth information of blur kernel and noise level. This indicates the effectiveness of the proposed blind framework which is capable of handling the tasks of degradation estimation and image restoration simultaneously. Further, taking the model profiles into consideration, the superiority of VIRNet is more evident. Specifically, VIRNet has fewer number of parameters, fewer FLOPs, and faster speed than both the SotA blind method DASR [70] and the non-blind method USRNet [56].
@@
V-B2 Results on Real Data
@@
V-B3 Discussion on Blur Kernel
@@
V-C Degradation Estimation Experiments
@@ V-C1 Noise Estimation
Non-I.I.D. Gaussian Noise. In Sec. V-A1, we adopt three groups of noise variance maps (see Fig. 3 (b1-d1)) to synthesize the testing data, to evaluate the performance of VIRNet under the non-i.i.d. Gaussian noise. Correspondingly, Fig. 3 (b2-d2) further displays the variance maps predicted by VIRNet for easy visualization. It can be seen that these predicted variance maps have very similar spatial variation with the ground truth ones, which are expected to facilitate the subsequent denoising task or other non-blind denoising methods. To justify this point, we also apply these predicted variance maps in FFDNet [20] to test its performance under the non-i.i.d. Gaussian noise, and the quantitative comparisons are listed in Table VIII. One can see that $\text{FFDNet}_{\text{VIR}}$ and $\text{FFDNet}_{\text{GT}}$ have very similar performance, and the performance difference between them is less than 0.04dB PSNR. This indicates that VIRNet is able to effectively handle such complicated noise distribution.
@@ V-C2 Kernel Estimation
V-D Ablation Study
V-D1 Posteriori Factorization
When designing the variational inference algorithm in Sec. IV, we factorize the variational distribution $q(\bm{z},\bm{\sigma}^{2},\bm{\varLambda}|\bm{y})$ into a conditional format of Eq. (12), which fundamentally induces the cascaded inference framework in Fig. 2. In fact, different factorized assumptions on $q(\bm{z},\bm{\sigma}^{2},\bm{\varLambda}|\bm{y})$ will lead to different designs on the inference framework. For example, the following unconditional factorization
@@ V-D1 Posteriori Factorization
V-D2 Hyper-parameter analysis
@@
Hyper-parameter $\varepsilon_{0}^{2}$. As discussed in Sec. IV-B2, Our ELBO loss degenerates to MSE when $\varepsilon_{0}$ is set as an extremely small value. We thus trained the RNet separately under MSE as a baseline for explicit comparison. The evaluative performance of VIRNet under different configurations of $\varepsilon_{0}$ is summarized in Table XII. This table unveils the following observations: 1) Too large setting on $\varepsilon_{0}^{2}$ (e.g., 1e-4) yields a relatively inferior performance of VIRNet, attributed to the attenuation of the prior constraint on $\bm{z}$; 2) our VIRNet performs well as long as $\varepsilon_{0}^{2}$ resides within the range of [1e-5, 1e-7], and deteriorates obviously when $\varepsilon_{0}$ is smaller beyond the threshold of 1e-7; 3) VIRNet surpasses the MSE baseline by 0.12dB in terms of PSNR, indicating the pivotal role of noise modeling in our proposed methodology. Thus, we suggest to set $\varepsilon_{0}^{2}$ as 1e-6 in this work.